### v2

# V. I. Telnov

Abstract
Recently, a high energy superconducting (SC) $e^{+}e^{-}$ linear collider (LC) with energy recovery (ERLC) has been proposed where twin RF structures are used to avoid parasitic collisions in linacs. Such a collider could operate in a DC (duty cycle) or CW (continuous) modes, if sufficient power is available, with a luminosity of ${\cal O}(10^{36})$ cm-2s-1 at $5002E_{0}=250-$ 500 GeV. In this paper, it is noted that the luminosity in the ERLC is independent of the accelerating gradient (at the same total power). Therefore, the ERLC could operate at maximum available accelerating gradients. This paper considers for the first time an $e^{-}e^{-}$ twin collider with ER and estimates its achievable luminosity. Such an $e^{-}e^{-}$ collider is much simpler than an $e^{+}e^{-}$ one because beam recirculation is not needed. Travelling wave RF structures allow for higher gradients and reduced heat load. It is shown that an ERLC with accelerating gradient of $G=40$ MeV/m could operate in CW mode with $L_{e^{+}e^{-}}=$ (1 – 2.5)×$\times 10^{36}$ and $L_{e^{-}e^{-}}=$ (3 – 7)×$\times 10^{36}$ cm-2s-1 for reasonable powers of 150 and 300 MW for $2502E_{0}=$ 250 and $5002E_{0}=$ 500 GeV, respectively. Such a machine is a promising candidate for a Higgs factory.
1 Introduction
Linear $e^{+}e^{-}$ colliders (LC) have been actively developed since the 1970s, as a way to achive higher energies. There were several LC projects in the 1990s (VLEPP, NLC, JLC, CLIC and TESLA, etc.), but only two remain today: ILC [1, 2] and CLIC [3]. The ILC uses superconducting (SC) Nb technology (in the footsteps of the TESLA), while the CLIC uses Cu cavities and operates at room temperature. Both colliders work in a pulsed mode and send beams to a beamdump after collision. The ILC has been ready for construction for many years, but no decision has been made due to uncertainty about the next collider following the LHC. At the same time, there are new ideas emerging on how to make linear colliders smaller, cheaper, or more luminosity. A recent review of all these approaches was prepared for the Snowmass-2021 conference and is available in ref. [4].
In this article, I continue to discuss my recent proposal for a twin superconducting $e^{+}e^{-}$ linear collider with energy recovery (ERLC) [5]. Its scheme is shown in Fig. 1. In this collider, beams are accelerated and then decelerated separately in parallel linacs with coupled RF systems. This prevents parasitic collisions that could destroy the beams. The same beams, e+ and e-, are used many times (>$10^{5}$), so all the advantages of the superconductivity are utilized. The achievable $e^{+}e^{-}$ luminosity ${\cal O}(10^{36})$ ${\rm\,cm}^{-2}{\rm s}^{-1}$ is much higher than that of a single-pass ILC. In my earlier publication [5], a relatively low accelerating gradient of $G=20$ MeV/m was assumed. At this gradient, the quality factor $Q$ of the SC cavities is close to its maximum. It may seem like a drawback, as the ILC plans to use an accelerating gradient around $G\approx 35$ MeV/m. Moreover, it has recently been noted that it is possible to nearly double the gradient by using a traveling wave (TW) instead of a standing wave (SW) [6]. This allows for gradients of 70 (Nd)–100 (Nb3Sn) MeV/m, although this technology is still under development.
In this article, I would like to draw your attention to the fact that the ERLC collider also works at high gradients. The luminosity does not depend on the acceleration gradient (at the same overall power), but it only slightly changes because of the dependence of the quality factor on gradient: $L\propto\sqrt{Q}$. Also, we consider for the first time the case of $e^{-}e^{-}$ ERLC. Such a $e^{-}e^{-}$ collider is much simpler than $e^{+}e^{-}$ one because beam recirculation isn’t required and a higher luminosity can be achieved. It is shown that ERLC $e^{+}e^{-}$ and $e^{-}e^{-}$ colliders can operate in continuous mode at $2E_{0}=$ 250–500 GeV and have luminosities higher than $10^{36}$ ${\rm\,cm}^{-2}{\rm s}^{-1}$ when using Nb3Sn and TW RF structures.
Figure 1: The layout of the SC twin $e^{+}e^{-}$ linear collider.
2 $e^{+}e^{-}$ ERLC: dependence of the luminosity on G and Q
Here we simply follow Ref. [5], where the required formula was derived and the dependence on $Q$ was emphasized, but no mention was made of the dependence on accelerating gradient.
Firstly, let’s consider assume the case where the collider operates with a duty cycle (DC), running for some time $DC<1$. This mode is possible at any average power level and at any high accelerating gradients. Let’s determine the optimal number of particles per bunch $N$ when the luminosity reaches its maximum for at a given power level.
There are two major energy consumers
  • Electric power to cool RF losses in low-temperature cavities, which does not depend on the particle number in the bunch.
  • Electric power required to compensate and remove of High Order Mode (HOM) losses. The HOM energy loss per unit length of a bunch is proportional to $N^{2}$. If the spacing between bunches $d$, then for the given collider $P_{\rm HOM}\propto N^{2}/d$.
The total power (only main contributions)
$P_{tot}=\left(k_{1}+k_{2}\frac{N^{2}}{d}\right)\times DC,$ (2.1)
where coefficients $k_{1}$ and $k_{2}$ (’$a$’ and ’$b$’ in [5]) describe RF and HOM losses, respectively, they both are proportional to the length of the collider (or $E_{0}$).
The $e^{+}e^{-}$ luminosity in one bunch collision $L_{1}$ is determined by the collision effects (beamstrahlung and bunch instability). For flat beams, these effects are the same when $N$ varies proportional to the horizontal beam size $\sigma_{x}$, therefore $L_{1}\propto N$, and the total luminosity
$L\propto\frac{N}{d}DC=\frac{N}{d}\left(\frac{P}{k_{1}+k_{2}N^{2}/d}\right).$ (2.2)
The maximum luminosity $L$
$L\propto\frac{P}{2\sqrt{k_{1}k_{2}d}}\;\;\;\;{\rm at}\;\;\;\;N=\sqrt{\frac{k_{1}d}{k_{2}}},\;\;\;\;DC=\frac{P}{2k_{1}}.$ (2.3)
The luminosity reaches its maximum when the energies spent on removing $RF$ and $HOM$ losses are equal. We see from (2.3) that for a fixed total power
  1. 1.
    $L\propto 1/\sqrt{d}$, so the distance between bunches $d$ should be as small as possible ($d=\lambda_{\rm RF}$ is the best);
  2. 2.
    $L\propto\sqrt{Q}$, because $k_{1}\propto 1/Q$;
  3. 3.
    $L$ does not depend on the acceleration gradient $G$. This is because the collider length $l_{c}\propto 1/G$, so $k_{2}\propto 1/G$, and the RF losses $k_{1}\propto G^{2}\times l_{c}\propto G$. As a result $L=\rm const$;
  4. 4.
    The optimal $N=\sqrt{k_{1}d/k_{2}}\propto G\sqrt{d/Q}$, $DC\propto 1/G$
We see, that the path towards a high ERLC accelerating gradient is open. The luminosity is independent of $G$, but weakly dependent on $Q$. This unexpected dependency is due to the optimal change in $N$ and $DC$. At $G=20$ MeV/m the optimum around $N\sim 10^{9}$ [5], so there is no difficulty in increasing it several times to achieve higher gradients.
In continuous (CW) mode
$P=(k_{1}+k_{2}\frac{N^{2}}{d}),\;\;\;L\propto\frac{N}{d},$ (2.4)
that gives
$N=\sqrt{\frac{(P-k_{1})d}{k_{2}}},\;\;\;\;\;L\propto\sqrt{\frac{(P-k_{1})}{k_{2}d}}.$ (2.5)
Continuous operation requires a large threshold power $P>k_{1}$. However, for higher powers, the behavior is similar to that of DC mode, where $L/P$ is independent of $G$ and is proportional to $Q^{1/2}$.
3 $e^{-}e^{-}$ ERLC
Let us consider a twin $e^{-}e^{-}$ linear collider with energy recovery. Such a collider is also of a great interest. Its physics program has been discussed at several dedicated workshops, and this option has always been taken into account when considering LC projects. A schematic diagram of such $e^{-}e^{-}$ collider can be seen in Fig. 2.
Figure 2: The layout of the SC twin $e^{-}e^{-}$ linear collider.
It is much simpler than a $e^{+}e^{-}$ collider, because no beam recycling is required. Electron beams with small emittances can be prepared each time anew. Beams are more tightly focussed than in a $e^{+}e^{-}$ case, since they are used only once. The difference between the ILC and ERLC is that in the latter, beams after collision return their energy to the RF field. Not all of it is recovered, some energy is lost through beamstrahlung at IP (interaction point) and must be replaced by the RF-system without energy recovery. In the case of $e^{+}e^{-}$ such average energy losses are insignificant. This requires extra electricity. In addition to average losses, there’s also a wide beam energy spread with tails. To fully utilize the beam’s energy, a decompressor can be placed at the end of the linac to reduce energy spread or several mini-beamdumps can be set up to remove low energy particles. For simplicity, we’ll assume twice the average energy loss needs to be compensated, with an efficiency $\varepsilon_{\rm rf}\sim 50$. Let’s consider it again, like we did for $e^{+}e^{-}$, but with an another source of energy consumption – beamstrahlung at the IP.
The relative energy loss due to beamstrahlung is give by [7]
$\delta=\frac{\Delta E}{E_{0}}\approx\frac{5r_{e}^{3}N^{2}\gamma}{6\sigma_{z}\sigma_{x}^{2}},$ (3.1)
where $r_{e}=e^{2}/mc^{2}$. The power required to compensate for these energy losses in two beams, multiplied by a factor of 2, as mentioned earlier, is
$P_{rad}=\frac{4NcE_{0}\delta}{\varepsilon_{\rm rf}d}=k_{3}\frac{N^{3}}{\sigma_{z}\sigma_{x}^{2}d},\,\,\,\,\,\,\,k_{3}=\frac{10E_{0}r_{e}^{3}\gamma c}{3\varepsilon_{\rm rf}}.$ (3.2)
3.1 $e^{-}e^{-}$ in duty cycle (DC) mode
Similar to (2.1), in $e^{-}e^{-}$ mode, the total power has three main components:
$P_{tot}=\left(k_{1}+k_{2}\frac{N^{2}}{d}+k_{3}\frac{N^{3}}{\sigma_{z}\sigma_{x}^{2}d}\right)\times DC.$ (3.3)
The luminosity
$L\approx\frac{N^{2}cH_{D}}{4\pi\sigma_{x}\sigma_{y}d}\times DC,$ (3.4)
where the vertical beam size $\sigma_{y}\approx\sqrt{\varepsilon_{ny}\sigma_{z}/\gamma}$, $H_{D}\sim 0.8$ is a geometric factor. Substituting (3.3) into (3.4) we get
$L=k_{0}\frac{N^{2}}{\sigma_{x}\sigma_{z}^{1/2}}\frac{P}{\left(k_{1}d+k_{2}N^{2}+k_{3}N^{3}/\sigma_{z}\sigma_{x}^{2}\right)},\,\,\,\,k_{0}=\frac{c\gamma^{1/2}H_{D}}{4\pi\varepsilon_{ny}^{1/2}}.$ (3.5)
The luminosity depends on two parameters: $N$ and $y=\sigma_{x}\sigma_{z}^{1/2}$
$L\propto\frac{N^{2}}{y\left(k_{1}d+k_{2}N^{2}+k_{3}N^{3}/y^{2}\right)}.$ (3.6)
For a fixed $N$ the maximum luminosity occurs when
$y^{2}=\frac{N^{3}k_{3}}{k_{1}d+k_{2}N^{2}}.$ (3.7)
At this value, the power consumption to compensate for radiation losses is 1/2 of the total power. Substituting (3.11) into (3.6) we get (only terms with $N$)
$L^{2}\propto\frac{NP^{2}}{k_{1}d+k_{2}N^{2}}.$ (3.8)
This dependance of $L^{2}$ on $N$ is the same as that of $L$ in the $e^{+}e^{-}$ collisions (2.2). The maximum luminosity occurs at the same value of $N$ value as in the $e^{+}e^{-}$ case.
$N=\sqrt{k_{1}d/k_{2}}.$ (3.9)
This implies that under optimal conditions $P_{\rm RF}=P_{\rm HOM}=0.5P_{rad}=0.25P$. The corresponding optimal duty cycle is
$DC=P/4k_{1},$ (3.10)
which is twice less than in the $e^{+}e^{-}$ case. The optimal value of $y^{2}$, describing beam sizes,
$y^{2}=\sigma_{x}^{2}\sigma_{z}=\frac{k_{1}^{1/2}k_{3}d^{1/2}}{2k_{2}^{3/2}}.$ (3.11)
Substituting all to (3.4) we obtain the maximum luminosity
$L=0.0154\frac{(c\varepsilon_{\rm rf})^{1/2}H_{D}P}{\varepsilon_{ny}^{1/2}(k_{1}k_{2}d)^{1/4}E_{0}^{1/2}r_{e}^{3/2}}.$ (3.12)
Above we obtained optimal values for $N$ and $\sigma_{x}^{2}\sigma_{z}$. In order to obtain separately $\sigma_{x}$ and $\sigma_{z}$ one should take into account the beam repulsion. It is determined by disruption parameter [7]
$D_{y}=\frac{2Nr_{e}\sigma_{z}}{\gamma\sigma_{x}\sigma_{y}}.$ (3.13)
For $e^{-}e^{-}$ collisions, the maximum value for $D_{y}\approx 5$ [8]. It does not affect the luminosity calculated above, but it determines the maximum bunch length. Substituting (3.9), (3.11) into (3.13) gives maximum value of $\sigma_{z}$ and corresponding value for $\sigma_{x}$.
[…]
3.2 $e^{-}e^{-}$ in continuous (CW) mode
Here we should repeat the considerations of the previous section, but for $DC=1$. In this case
$P_{tot}=\left(k_{1}+k_{2}\frac{N^{2}}{d}+k_{3}\frac{N^{3}}{\sigma_{z}\sigma_{x}^{2}d}\right)$ (3.14)
and
$L\approx\frac{N^{2}cH_{D}}{4\pi\sigma_{x}\sigma_{y}d}=\frac{k_{0}N^{2}H_{D}}{d\sigma_{x}\sigma_{z}^{1/2}},\,\,\,\,\;\;\;k_{0}=\frac{c\gamma^{1/2}H_{D}}{4\pi\varepsilon_{ny}^{1/2}},\;\;\;\;\;\sigma_{y}\approx\sqrt{\varepsilon_{ny}\sigma_{z}/\gamma}.$ (3.15)
Here, again, we have two parameters: $N$ and $\sigma_{x}^{2}\sigma_{z}$. The maximum luminosity occurs at
$N^{2}=\frac{(P-k_{1})d}{3k_{2}};\;\;\;\;\mbox{and}\;\;\;\;\;\sigma_{x}^{2}\sigma_{z}=\frac{k_{3}(P-k_{1})^{1/2}d^{1/2}}{2\sqrt{3}\,k_{2}^{3/2}}.$ (3.16)
In CW mode, the total power $P$ is distributed among RF, HOM and beamstrahlung losses as follows:
$k_{1}\;\;\;:\;\;\;\frac{P-k_{1}}{3}\;\;\;:\;\;\;\frac{2}{3}(P-k_{1}).$ (3.17)
The optimized luminosity in the CW mode is given by
$L=0.027\frac{(c\varepsilon_{\rm rf})^{1/2}H_{D}(P-k_{1})^{3/4}}{\varepsilon_{ny}^{1/2}(k_{2}d)^{1/4}E_{0}^{1/2}r_{e}^{3/2}}.$ (3.18)
Continuous operation requires a large threshold power $P>k_{1}$. For higher powers, the behavior is similar to that of DC mode, where $L/P$ does not depend on $G$ and is proportional to $Q^{1/4}$.
4 Examples
Let’s move on to the luminosity estimates, using numbers for $k_{1}$ and $k_{2}$ from ref. [5]. For Nd ILC-like cavities with $f_{\rm rf}=1.3$ GHz and $Q=3\times 10^{1}0$, the values $k_{1}\approx 305$ MW, $k_{2}\approx(240\,{\rm MW}/10^{18})\times 23{\rm cm}$ for $2502E_{0}=$ 250 GeV and $G=20$ MeV/m. For Nb3Sn cavities the value of $k_{1}$ is 4 times lower due to higher cryogenic efficiency. In the case of BCS surface conductivity $k_{1}\propto f_{\rm rf}$, $k_{2}\propto f_{\rm rf}^{2}$, $d\propto 1/f_{\rm rf}$. As discussed earlier, $k_{1}\propto E_{0}G$ and $k_{2}\propto E_{0}/G$.
Additional improvement can be achieved by using traveling wave (TW) structures instead of standing wave cavities used at the ILC [6, 9]. So, using TW structures with an aperture radius of 3.5 cm (as in the ILC) one can achieve 1.3 times greater G and 1.7 times lower RF heat loss (parameter $k_{1}$). Such a reduction in RF losses is very important because it allows for CW operation of an ERLC collider at $2502E=$ 250 and 500 GeV with $G=40$ MeV/m, at an acceptable total power.
The ERLC collider with a very high luminosity can be considered as a candidate for a future collider following the ILC in the same tunnel [10]. In this scenario, the acceleration gradient must be at least as high as that of the ILC, let’s assume $G=40$ MeV/m. The most promising (but not yet available) technologies are Nb3Sn ($f=1.3$ or 0.65 GHz) and traveling wave RF structures. It is advantageous to operate in CW mode as it is better for cooling systems. Based on these assumption and future vision, we present below igures and tables for possible parameters for $e^{+}e^{-}$ and $e^{-}e^{-}$ colliders at energies of $2E=$ 250 and 500 GeV.
Figure 3: Dependence of the $e^{+}e^{-}$ luminosity on the total power for $2502E_{0}=$ 250 GeV (upper) and $5002E_{0}=$ 500 GeV (bottom) in continuous mode of operation (CW), see the text.
Figure 4: Dependence of the $e^{-}e^{-}$ luminosity on the total power for $2502E_{0}=$ 250 GeV (upper) and $5002E_{0}=$ 500 GeV (bottom) in continuous(CW) and duty cycle(DC) modes of operation, see the text.
Table 1: Parameters of $e^{+}e^{-}$ and $e^{-}e^{-}$ linear colliders ERLC (Nb3Sn), $2502E_{0}=$ 250 GeV
unit $e^{+}e^{-}$ $e^{+}e^{-}$ $e^{-}e^{-}$ $e^{-}e^{-}$
1.3 GHz 0.65 GHz 1.3 GHz 0.65 GHz
Energy $2E_{0}$ GeV 250 250 250 250
Luminosity ${\mathcal{L}}_{{\rm tot}}$ $10^{36}$${\rm\,cm}^{-2}{\rm s}^{-1}$ 1.1 2.5 3.8 7.8
$P$ (wall) (collider) MW 150 150 150 150
Accel. gradient, $G$ MV/m 40 40 40 40
Length $L_{\rm act}/L_{\rm tot}$ km 6.3/16 6.3/16 6.3/16 6.3/16
$N$ per bunch $10^{9}$ 0.62 2.75 0.41 1.53
Bunch distance m 0.23 0.46 0.23 0.46
Rep. rate, $f$ Hz $1.3\cdot 10^{9}$ $0.65\cdot 10^{9}$ $1.3\cdot 10^{9}$ $0.65\cdot 10^{9}$
$\epsilon_{n,\,x}$/$\epsilon_{n,\,y}$ $10^{-6}$ m 10/0.03 10/0.03 1/0.02 1/0.02
$\beta^{*}_{x}$/$\beta_{y}$ at IP cm 0.82/0.03 16.4/0.03 0.21/0.019 3/0.02
$\sigma_{x}$ at IP ${\,\mu\rm m}$ 0.58 2.6 0.092 0.35
$\sigma_{y}$ at IP nm 6.2 6.2 4 4.05
$\sigma_{z}$ at IP cm 0.031 0.031 0.019 0.02
Table 2: Parameters of $e^{+}e^{-}$ and $e^{-}e^{-}$ linear colliders ERLC (Nb3Sn), $5002E_{0}=$ 500 GeV
unit $e^{+}e^{-}$ $e^{+}e^{-}$ $e^{-}e^{-}$ $e^{-}e^{-}$
1.3 GHz 0.65 GHz 1.3 GHz 0.65 GHz
Energy $2E_{0}$ GeV 500 500 500 500
Luminosity ${\mathcal{L}}_{{\rm tot}}$ $10^{36}$${\rm\,cm}^{-2}{\rm s}^{-1}$ 1.0 1.9 3.8 6.9
$P$ (wall) (collider) MW 300 300 300 300
Accel. gradient, $G$ MV/m 40 40 40 40
Length $L_{\rm act}/L_{\rm tot}$ km 12.5/30 12.5/30 12.5/30 12.5/30
$N$ per bunch $10^{9}$ 0.81 2.97 0.41 1.53
Bunch distance m 0.23 0.46 0.23 0.46
Rep. rate, $f$ Hz $1.3\cdot 10^{9}$ $0.65\cdot 10^{9}$ $1.3\cdot 10^{9}$ $0.65\cdot 10^{9}$
$\epsilon_{n,\,x}$/$\epsilon_{n,\,y}$ $10^{-6}$ m 10/0.03 10/0.03 1/0.02 1/0.02
$\beta^{*}_{x}$/$\beta_{y}$ at IP cm 4/0.09 55/0.09 0.41/0.039 6/0.04
$\sigma_{x}$ at IP ${\,\mu\rm m}$ 0.91 3.4 0.092 0.35
$\sigma_{y}$ at IP nm 7.4 7.4 4 4.05
$\sigma_{z}$ at IP cm 0.09 0.09 0.039 0.04
5 Discussion
As one can see, the attainable $L_{e^{+}e^{-}}=(1-2.5)10^{36}$ and $L_{e^{-}e^{-}}=(3-7)10^{36}$ ${\rm\,cm}^{-2}{\rm s}^{-1}$ for reasonable powers of 150 MW and 300 MW for $2502E_{0}=$ 250 and $5002E_{0}=$ 500 GeV, respectively. Cavities at 0.65 GHz allow operation in CW mode at about 2 times lower total power, giving about 2×2$\times$2 × higher luminosity, although they require wider tunnels.
An electron-positron collider is more complicated, as the same bunches must be used many times. This required measures to preserve beam emittance. In fact, it’s like a storage ring with additional components inside, such as linacs and bunch compressors.
Electron-electron ERLC collider is simpler, as each bunch is used only once. Their luminosity is very high, up to $10^{37}$ ${\rm\,cm}^{-2}{\rm s}^{-1}$. Their physics program is also very reach. It’s an attractive option.
6 Conclusion
Twin $e^{+}e^{-}$ and $e^{-}e^{-}$ linear colliders with the energy recovery open the way to extremely high luminosities. This article demonstrates that the luminosity of the ERLC colliders does not depend on the accelerating gradient and only weakly depends on quality factor of accelerating cavities: $L_{e^{+}e^{-}}\propto Q^{1/2}$, $L_{e^{-}e^{-}}\propto Q^{1/4}$. In the previous paper [5], $e^{+}e^{-}$ ERLC with repetitive use of bunches were considered; this paper considers the case of $e^{-}e^{-}$ ERLC for the first time with single use of electron bunches. The achivable $L_{e^{+}e^{-}}=$ ( 1 – 2.5)×$\times 10^{36}$ for $2E_{0}=$ 250 – 500 GeV and about 3 times higher for $e^{-}e^{-}$ collisions. Such a collider, with energy recovery, can be considered the next (high luminosity) stage of ILC (in its tunnel). Energy recovery superconducting accelerators have many potential applications [11], so we can expect rapid progress in this field.
Acknowledgment
This work was supported by the Russian Science Foundation with grant number 24-22-00288.