### v3

# Energy recovery twin linear $e^{+}e^{-}$, $e^{-}e^{-}$ colliders (ERLC ) with high luminosities and acceleration gradients

Abstract
Recently, a high-energy superconducting $e^{+}e^{-}$ linear collider (LC) with energy recovery (ERLC) has been proposed, using twin RF structures to avoid parasitic collisions in linacs. Such a collider could operate in DC (duty cycle) or CW (continuous) mode, if sufficient power is available, with a luminosity of ${\cal O}(10^{36})$ cm-2s-1 at $2E_{0}=$250 – 500 GeV. In this paper, we find that the luminosity of the ERLC operating in DC mode does not depend on the accelerating gradient (for the same total power), allowing the ERLC to operate at the maximum available accelerating gradients. We also consider for the first time an $e^{-}e^{-}$ twin collider with energy recovery and estimate its achievable luminosity. Such an $e^{-}e^{-}$ collider is simpler than an $e^{+}e^{-}$ one. It does not require beam recycling, since electrons can be produced anew each time. Travelling-wave RF structures allow higher gradients and reduced thermal loading. It is shown that an ERLC with an acceleration gradient of $G=40$ MeV/m can be operated in CW mode with $L_{e^{+}e^{-}}=$ (1 – 2.5) $10^{36}$ and $L_{e^{-}e^{-}}=$ (3 – 7) $10^{36}$ cm-2s-1 for reasonable powers of 150 and 300 MW at $2502E_{0}=$ 250 and 500 GeV, respectively. Such a machine is a promising candidate for a Higgs factory.
1 Introduction
Linear $e^{+}e^{-}$ colliders (LC) have been actively developed since the 1970s, as a way to achieve higher energies. There were several LC projects in the 1990s (VLEPP, NLC, JLC, CLIC and TESLA, etc.), but only two remain today: ILC [1, 2] and CLIC [3]. The ILC uses superconducting (SC) Nb technology (in the footsteps of the TESLA), while the CLIC uses Cu cavities and operates at room temperature. Both colliders work in a pulsed mode and send beams to a beamdump after collision. The ILC has been ready for construction for many years, but no decision has been made due to uncertainty about the next collider following the LHC. At the same time, there are new ideas emerging on how to make linear colliders smaller, cheaper, or more luminosity. A recent review of all these approaches was prepared for the Snowmass-2021 conference and is available in ref. [4].
In this article I continue to discuss my recent proposal for a twin superconducting $e^{+}e^{-}$ linear collider with energy recovery (ERLC) [5]. Its scheme is shown in Fig. 1. In this collider, beams are accelerated and then decelerated separately in parallel linacs with coupled RF systems. This prevents parasitic collisions that could destroy the beams. The same beams, e+ and e-, are used many times (>$10^{5}$), so that all the advantages of superconductivity are exploited. The achievable $e^{+}e^{-}$ luminosity ${\cal O}(10^{36})$ ${\rm\,cm}^{-2}{\rm s}^{-1}$ is much higher than that of a single-pass ILC. In my earlier publication [5], a relatively low accelerating gradient of $G=20$ MeV/m was assumed. This may seem like a disadvantage, since the ILC plans to use a gradient of around 35 MeV/m. In addition, it was recently discovered that it is possible to increase the gradient even further by using a traveling wave (TW) instead of a standing wave (SW) [6]. This allows gradients of 70 (Nd)–100 (Nb3Sn) MeV/m, although this technology is still under development.
This article shows that the ERLC collider can also work at high gradients. In DC mode, the luminosity does not depend on the acceleration gradient (for the same total power), and it changes only slightly because of the dependence of a cavity quality factor Q on the acceleration gradient G. We also consider for the first time the case of $e^{-}e^{-}$ ERLC. Such a $e^{-}e^{-}$ collider is much simpler than a $e^{+}e^{-}$ one, because it doesn’t require beam recirculation and can achieve higher luminosity. Finally, we show that ERLC $e^{+}e^{-}$ and $e^{-}e^{-}$ colliders can operate in continuous mode (CW) at $2E_{0}=$ 250–500 GeV and have luminosities higher than $10^{36}$ ${\rm\,cm}^{-2}{\rm s}^{-1}$ using Nb3Sn and TW RF structures.
Figure 1: The layout of the SC twin $e^{+}e^{-}$ linear collider [5]
2 ERLC-$e^{+}e^{-}$: dependence of L on G and Q
Here we just follow Ref. [5], where all the necessary formulas were derived, but the dependence on the acceleration gradient was not explicitly emphasized.
First, let’s consider the case where the collider operates with a duty cycle (DC). This mode is possible at any average power level and at any high accelerating gradients. There are two main consumers of energy:
  • Electrical power to cool the RF losses in the low-temperature cavities, which is independent of the number of particles in the bunch,
  • Electrical power required to compensate for and remove High Order Mode (HOM) losses. The HOM energy loss per unit length of a bunch is proportional to $N^{2}$. If the distance between the bunches is $d$, then for the given collider $P_{\rm HOM}\propto N^{2}/d$.
The total power (main contributions only)
$P_{tot}=\left(k_{1}+k_{2}\frac{N^{2}}{d}\right)\times DC,$ (2.1)
where the coefficients $k_{1}$ and $k_{2}$ (’$a$’ and ’$b$’ in [5]) describe the RF and HOM losses, respectively, both of which are proportional to the length of the collider (or $E_{0}$).
The $e^{+}e^{-}$ luminosity in a bunch collision $L_{1}$ is determined by the collision effects (beam radiation and bunch instability). For flat beams, these effects are the same if $N$ varies proportionally to the horizontal beam size $\sigma_{x}$, hence $L_{1}\propto N$, and the total luminosity is
$L\propto\frac{N}{d}DC=\frac{N}{d}\left(\frac{P}{k_{1}+k_{2}N^{2}/d}\right).$ (2.2)
The maximum luminosity
$L\propto\frac{P}{2\sqrt{k_{1}k_{2}d}}\;\;\;\;\;{\rm at}\;\;\;\;\;N=\sqrt{\frac{k_{1}d}{k_{2}}},\;\;\;\;\;\;DC=\frac{P}{2k_{1}}.$ (2.3)
Luminosity reaches a maximum when the energy consumption for RF and HOM losses are equal. We see from (2.3) that for a fixed total power
  1. 1.
    $L\propto 1/\sqrt{d}$, so the distance between the bunches $d$ should be as small as possible ($d=\lambda_{\rm RF}$ is the best);
  2. 2.
    $L\propto\sqrt{Q}$, because $k_{1}\propto 1/Q$;
  3. 3.
    $L$ does not depend on the acceleration gradient $G$. This is because the collider length is $l_{c}\propto 1/G$, so $k_{2}\propto 1/G$, and the RF losses are $k_{1}\propto G^{2}\times l_{c}\propto G$. The result is $L=\rm const$;
  4. 4.
    The optimal $N=\sqrt{k_{1}d/k_{2}}\propto G\sqrt{d/Q}$, $DC\propto 1/G$.
We see that the path to a high ERLC acceleration gradient is open. The luminosity is independent of $G$, but weakly dependent on $Q$. This unexpected dependence is due to the optimal variation of $N$ and $DC$. At $G=20$ MeV/m the optimum is around $N\sim 10^{9}$ [5], so there is no difficulty to increase it several times to get higher gradients.
In continuous mode (CW)
$P=(k_{1}+k_{2}\frac{N^{2}}{d}),\;\;\;L\propto\frac{N}{d},$ (2.4)
which gives
$N=\sqrt{\frac{(P-k_{1})d}{k_{2}}},\;\;\;\;\;\;L\propto\sqrt{\frac{(P-k_{1})}{k_{2}d}}.$ (2.5)
Continuous operation requires a large threshold power $P>k_{1}\propto G$. However, at higher powers the behaviour is similar to DC mode, where $L/P$ does not depend on $G$ and proportional to $Q^{1/2}$.
3 ERLC-$e^{-}e^{-}$
Let us consider a twin $e^{-}e^{-}$ linear collider with energy recovery. Such a collider is also of a great interest. Its physics program has been discussed in several dedicated workshops, and this option has always been taken into account when considering LC projects. A schematic diagram of such a $e^{-}e^{-}$ collider can be seen in Fig. 2.
Figure 2: The layout of the SC twin $e^{-}e^{-}$ linear collider.
It is much simpler than a $e^{+}e^{-}$ collider because no beam recycling is required. Low emittance electron beams can be prepared anew each time (without damping rings). The beams can be more tightly focused (and more disrupted) than in a $e^{+}e^{-}$ because they are used only once. The difference between the ILC and the ERLC is that in the latter, the beams return their energy to the RF field after the collision. Not all of it is recovered, some energy is lost through beam radiation at the IP (interaction point) and must be replaced by the RF system without energy recovery. This requires additional power. In the case of $e^{+}e^{-}$, such average energy losses are insignificant. In addition to average losses, there’s also a wide beam energy spread with tails. To fully utilize the beam energy, a decompressor must be placed at the end of the linac to reduce the energy spread. The lowest-energy particles can be removed using several mini-beam dumps. For simplicity, we’ll assume that twice the average energy loss needs to be compensated, with an efficiency of $\varepsilon_{\rm rf}\sim 50$.
Let’s calculate the $e^{-}e^{-}$ luminosity taking into account an additional source of energy consumption – the beam radiation (beamstrahlung) at the interaction point (IP).
The relative energy loss due to beamstrahlung is given by [7]
$\delta=\frac{\Delta E}{E_{0}}\approx\frac{5r_{e}^{3}N^{2}\gamma}{6\sigma_{z}\sigma_{x}^{2}},$ (3.1)
where $r_{e}=e^{2}/mc^{2}$. The power required to compensate for these energy losses in two beams, multiplied by a factor of 2, as mentioned above, is
$P_{rad}=\frac{4NcE_{0}\delta}{\varepsilon_{\rm rf}d}=k_{3}\frac{N^{3}}{\sigma_{z}\sigma_{x}^{2}d},\,\,\,\,\,\,\,k_{3}=\frac{10E_{0}r_{e}^{3}\gamma c}{3\varepsilon_{\rm rf}}.$ (3.2)
3.1 $e^{-}e^{-}$ DC mode
Unlike (2.1), the total power in $e^{-}e^{-}$ mode consists of three main components:
$P_{tot}=\left(k_{1}+k_{2}\frac{N^{2}}{d}+k_{3}\frac{N^{3}}{\sigma_{z}\sigma_{x}^{2}d}\right)\times DC.$ (3.3)
The $e^{-}e^{-}$ luminosity
$L\approx\frac{N^{2}cH_{D}}{4\pi\sigma_{x}\sigma_{y}d}\times DC,$ (3.4)
where the vertical beam size $\sigma_{y}\approx\sqrt{\varepsilon_{ny}\sigma_{z}/\gamma}$, $H_{D}\sim 0.8$ is a geometric factor. Substituting (3.3) into (3.4) we get
$L=k_{0}\frac{N^{2}}{\sigma_{x}\sigma_{z}^{1/2}}\frac{P}{\left(k_{1}d+k_{2}N^{2}+k_{3}N^{3}/\sigma_{z}\sigma_{x}^{2}\right)},\,\,\,\,k_{0}=\frac{c\gamma^{1/2}H_{D}}{4\pi\varepsilon_{ny}^{1/2}}.$ (3.5)
The luminosity depends on two parameters: $N$ and $y=\sigma_{x}\sigma_{z}^{1/2}$ and is proportional to
$L\propto\frac{N^{2}}{y\left(k_{1}d+k_{2}N^{2}+k_{3}N^{3}/y^{2}\right)}.$ (3.6)
For a fixed $N$ the maximum luminosity occurs when
$y^{2}=\frac{N^{3}k_{3}}{k_{1}d+k_{2}N^{2}}.$ (3.7)
At this value, the power required to compensate for radiation losses is 1/2 of the total power. Substituting (3.7) into (3.6) we get (only terms with $N$)
$L^{2}\propto\frac{NP^{2}}{k_{1}d+k_{2}N^{2}}.$ (3.8)
This dependence of $L^{2}$ on $N$ is the same as the dependence of $L$ on $e^{+}e^{-}$ collisions (2.2). The maximum luminosity occurs at the same value of $N$ value as in the $e^{+}e^{-}$ case.
$N=\sqrt{k_{1}d/k_{2}}.$ (3.9)
This implies that under optimal conditions $P_{\rm RF}=P_{\rm HOM}=0.5P_{rad}=0.25P$. The corresponding optimal duty cycle is
$DC=P/4k_{1},$ (3.10)
which is twice smaller than in the case of $e^{+}e^{-}$. For optimal $N$, the value of $y^{2}$, which describes beam sizes, is
$y^{2}=\sigma_{x}^{2}\sigma_{z}=\frac{k_{1}^{1/2}k_{3}d^{1/2}}{2k_{2}^{3/2}}.$ (3.11)
Substituting (3.9),(3.10) and (3.11) in (3.4) we get the maximum luminosity
$L=0.0154\frac{(c\varepsilon_{\rm rf})^{1/2}H_{D}P}{\varepsilon_{ny}^{1/2}(k_{1}k_{2}d)^{1/4}E_{0}^{1/2}r_{e}^{3/2}}.$ (3.12)
Above we have obtained optimal values for $N$ and $\sigma_{x}^{2}\sigma_{z}$. To obtain $\sigma_{x}$ and $\sigma_{z}$ separately, one should take into account the beam repulsion. It is determined by the disruption parameter [7]
$D_{y}=\frac{2Nr_{e}\sigma_{z}}{\gamma\sigma_{x}\sigma_{y}}.$ (3.13)
For $e^{-}e^{-}$ collisions, the maximum value for $D_{y}\approx 5$ [8]. It does not affect the luminosity calculated above, but it does determines the maximum bunch length. By substituting (3.9), (3.11) into (3.13) we get the maximum value of $\sigma_{z}$, and then using (3.11, we find the corresponding value of $\sigma_{x}$.
[…]
3.2 $e^{-}e^{-}$ CW mode
Here we need to repeat the considerations of the previous section, but for $DC=1$. In this case
$P_{tot}=\left(k_{1}+k_{2}\frac{N^{2}}{d}+k_{3}\frac{N^{3}}{\sigma_{z}\sigma_{x}^{2}d}\right)$ (3.14)
and
$L\approx\frac{N^{2}cH_{D}}{4\pi\sigma_{x}\sigma_{y}d}=\frac{k_{0}N^{2}H_{D}}{d\sigma_{x}\sigma_{z}^{1/2}},\,\,\,\,\;\;\;k_{0}=\frac{c\gamma^{1/2}H_{D}}{4\pi\varepsilon_{ny}^{1/2}},\;\;\;\;\;\sigma_{y}\approx\sqrt{\varepsilon_{ny}\sigma_{z}/\gamma}.$ (3.15)
Again, we have two parameters: $N$ and $\sigma_{x}^{2}\sigma_{z}$. The maximum luminosity is reached at
$N^{2}=\frac{(P-k_{1})d}{3k_{2}};\;\;\;\;\mbox{and}\;\;\;\;\;\sigma_{x}^{2}\sigma_{z}=\frac{k_{3}(P-k_{1})^{1/2}d^{1/2}}{2\sqrt{3}\,k_{2}^{3/2}}.$ (3.16)
In $e^{-}e^{-}$ CW mode, the total power $P$ is distributed among RF, HOM and beamstrahlung losses as follows:
$k_{1}\;\;\;:\;\;\;\frac{P-k_{1}}{3}\;\;\;:\;\;\;\frac{2}{3}(P-k_{1}).$ (3.17)
The optimized luminosity in the $e^{+}e^{-}$ CW mode is given by
$L=0.027\frac{(c\varepsilon_{\rm rf})^{1/2}H_{D}(P-k_{1})^{3/4}}{\varepsilon_{ny}^{1/2}(k_{2}d)^{1/4}E_{0}^{1/2}r_{e}^{3/2}}.$ (3.18)
Continuous operation requires a large threshold power $P>k_{1}\propto G$. For higher powers, the behavior is similar to that of DC mode, where $L/P$ does not depend on $G$ and is proportional to $Q^{1/4}$.
4 Examples
Let’s move on to the luminosity estimates, using the values for $k_{1}$ and $k_{2}$ from ref. [5]. For Nd ILC-like cavities with $f_{\rm rf}=1.3$ GHz and $Q=3\times 10^{10}$, the values $k_{1}\approx 305$ MW, $k_{2}\approx(240\,{\rm MW}/10^{18})\times 23{\rm cm}$ for $2502E_{0}=$ 250 GeV and $G=20$ MeV/m. For Nb3Sn cavities the value of $k_{1}$ is 4 times lower due to the higher cryogenic efficiency. In the case of BCS surface conductivity $k_{1}\propto f_{\rm rf}$, $k_{2}\propto f_{\rm rf}^{2}$, $d\propto 1/f_{\rm rf}$. As discussed earlier, $k_{1}\propto E_{0}G$ and $k_{2}\propto E_{0}/G$.
Important note: the value of $k_{2}$ for HOM losses given above was obtained under the assumption that the HOM loss of the bunch train is the sum of the losses of each bunch separately, which is valid only if the phases are random. In the ERLC, as noted in the same ref. [5], the bunches are strictly periodic, in this case a significant reduction of HOM losses can be expected. This issue is very important and needs to be studied in detail.
An additional improvement can be achieved by using traveling wave (TW) structures [6, 9] instead of the standing wave cavities used in the ILC. Thus, using TW cavities with an aperture radius of 3.5 cm (as in the ILC), one can achieve 1.3 times higher G and 1.7 times lower RF heat loss (parameter $k_{1}$[9]. Such a reduction in RF losses is very important because it allows for CW operation of an ERLC collider at $2502E=$ 250 and 500 GeV with a high acceleration gradient at an acceptable total power.
The ERLC collider with a very high luminosity can be considered as a candidate for a future collider following the ILC in the same tunnel [10]. In this scenario, the acceleration gradient must be at least as high as that of the ILC, let’s assume $G=40$ MeV/m. The most promising (but not yet available) technologies are Nb3Sn ($f=1.3$ or 0.65 GHz) and traveling wave RF structures. It is advantageous to operate in CW mode as it is better for cooling systems. Based on these assumptions and future visions, we present below figures and tables for possible parameters for $e^{+}e^{-}$ and $e^{-}e^{-}$ colliders at energies of $2E_{0}=$ 250 and 500 GeV.
Figure 3: Dependence of the $e^{+}e^{-}$ luminosity on the total power for $2502E_{0}=$ 250 GeV (upper) and $5002E_{0}=$ 500 GeV (bottom) in continuous mode of operation (CW), see the text.
Figure 4: Dependence of the $e^{-}e^{-}$ luminosity on the total power for $2502E_{0}=$ 250 GeV (upper) and $5002E_{0}=$ 500 GeV (bottom) in continuous(CW) and duty cycle(DC) modes of operation, see the text.
Table 1: Parameters of $e^{+}e^{-}$ and $e^{-}e^{-}$ linear colliders ERLC (Nb3Sn), $2502E_{0}=$ 250 GeV
unit $e^{+}e^{-}$ $e^{+}e^{-}$ $e^{-}e^{-}$ $e^{-}e^{-}$
1.3 GHz 0.65 GHz 1.3 GHz 0.65 GHz
Energy $2E_{0}$ GeV 250 250 250 250
Luminosity ${\mathcal{L}}_{{\rm tot}}$ $10^{36}$${\rm\,cm}^{-2}{\rm s}^{-1}$ 1.1 2.5 3.8 7.8
$P$ (wall) (collider) MW 150 150 150 150
Accel. gradient, $G$ MV/m 40 40 40 40
Length $L_{\rm act}/L_{\rm tot}$ km 6.3/16 6.3/16 6.3/16 6.3/16
$N$ per bunch $10^{9}$ 0.62 2.75 0.41 1.53
Bunch distance m 0.23 0.46 0.23 0.46
Rep. rate, $f$ Hz $1.3\cdot 10^{9}$ $0.65\cdot 10^{9}$ $1.3\cdot 10^{9}$ $0.65\cdot 10^{9}$
$\epsilon_{n,\,x}$/$\epsilon_{n,\,y}$ $10^{-6}$ m 10/0.03 10/0.03 1/0.02 1/0.02
$\beta^{*}_{x}$/$\beta_{y}$ at IP cm 0.82/0.03 16.4/0.03 0.21/0.019 3/0.02
$\sigma_{x}$ at IP ${\,\mu\rm m}$ 0.58 2.6 0.092 0.35
$\sigma_{y}$ at IP nm 6.2 6.2 4 4.05
$\sigma_{z}$ at IP cm 0.031 0.031 0.019 0.02
Table 2: Parameters of $e^{+}e^{-}$ and $e^{-}e^{-}$ linear colliders ERLC (Nb3Sn), $5002E_{0}=$ 500 GeV
unit $e^{+}e^{-}$ $e^{+}e^{-}$ $e^{-}e^{-}$ $e^{-}e^{-}$
1.3 GHz 0.65 GHz 1.3 GHz 0.65 GHz
Energy $2E_{0}$ GeV 500 500 500 500
Luminosity ${\mathcal{L}}_{{\rm tot}}$ $10^{36}$${\rm\,cm}^{-2}{\rm s}^{-1}$ 1.0 1.9 3.8 6.9
$P$ (wall) (collider) MW 300 300 300 300
Accel. gradient, $G$ MV/m 40 40 40 40
Length $L_{\rm act}/L_{\rm tot}$ km 12.5/30 12.5/30 12.5/30 12.5/30
$N$ per bunch $10^{9}$ 0.81 2.97 0.41 1.53
Bunch distance m 0.23 0.46 0.23 0.46
Rep. rate, $f$ Hz $1.3\cdot 10^{9}$ $0.65\cdot 10^{9}$ $1.3\cdot 10^{9}$ $0.65\cdot 10^{9}$
$\epsilon_{n,\,x}$/$\epsilon_{n,\,y}$ $10^{-6}$ m 10/0.03 10/0.03 1/0.02 1/0.02
$\beta^{*}_{x}$/$\beta_{y}$ at IP cm 4/0.09 55/0.09 0.41/0.039 6/0.04
$\sigma_{x}$ at IP ${\,\mu\rm m}$ 0.91 3.4 0.092 0.35
$\sigma_{y}$ at IP nm 7.4 7.4 4 4.05
$\sigma_{z}$ at IP cm 0.09 0.09 0.039 0.04
5 Discussion
As one can see, the achievable $L_{e^{+}e^{-}}$ = (1 – 2.5)$10^{36}$ and $L_{e^{-}e^{-}}$ = (3 – 7)$10^{36}$ ${\rm\,cm}^{-2}{\rm s}^{-1}$ for reasonable powers of 150 MW and 300 MW at $2502E_{0}=$ 250 and $5002E_{0}=$ 500 GeV, respectively. Compared to $f_{\rm RF}=1.3$ GHz, cavities operating at 0.65 GHz allow CW operation at about half the power and provide about twice the luminosity, although they require wider tunnels.
An electron-positron collider is more complicated because the same bunches have to be used many times. This requires measures to preserve the emittance of the beam. In fact, it is like a storage ring with additional components inside, such as linear accelerators and bunch compressors.
The ERLC electron-electron collider is simpler because each bunch is used only once (it can be prepared anew each time). Its luminosity can be very high, so it’s also an attractive option.
6 Conclusion
Twin $e^{+}e^{-}$ and $e^{-}e^{-}$ linear colliders with the energy recovery open the way to extremely high luminosities. This article shows that the luminosity of the ERLC colliders does not depend on the acceleration gradient and only weakly on the quality factor of the RF cavities: $L_{e^{+}e^{-}}\propto Q^{1/2}$, $L_{e^{-}e^{-}}\propto Q^{1/4}$. In the previous paper [5], $e^{+}e^{-}$ ERLC with repeated use of bunches was considered; this paper also considers for the first time the case of $e^{-}e^{-}$ ERLC with single use of electron bunches. The achievable $e^{+}e^{-}$ luminosity $L_{e^{+}e^{-}}=$ ( 1 – 2.5) $10^{36}$ for $2E_{0}=$ 250 – 500 GeV and about 3 times higher in $e^{-}e^{-}$ collisions. Such an energy-recovery collider can be considered as the next high-luminosity stage of the ILC (in its tunnel). Energy-recovery superconducting accelerators have many potential applications [11], so one can expect rapid progress in this field.
Acknowledgment
This work was supported by the Russian Science Foundation with grant number 24-22-00288.